# Density-valued time series: Nonparametric density-on-density regression

### v1

@@ 3.1 Implementation
Since the actual densities in practical applications are not often observable, we work with densities which are outputs of kernel density estimators:
$f_{t}(u)=\frac{1}{n_{t}h_{t}}\sum^{n_{t}}_{i=1}K\left(\frac{u-X_{t,i}}{h_{t}}\right),\qquad t=1,\dots,n,$ (3)
where $K(\cdot)$ denotes a kernel function and $h_{t}$ denotes a bandwidth for period $t$. We consider the truncated Gaussian kernel.
In our nonparametric density-on-density regression, we consider how to select the optimal bandwidths for estimating the nonparametric regression operator in (1) and for estimating densities in (3). To estimate densities from observed data, we use a kernel density estimator with the bandwidth selected by Silverman’s rule-of-thumb (ROT) (see also Kokoszka et al., 2019). Kokoszka et al. (2019) used Silverman’s ROT to select the bandwidth, which leads to $\widehat{h}_{t}=2.34\times\widehat{\sigma}_{t}\times n_{t}^{-1/5}$, where $\widehat{\sigma_{t}}$ denotes the sample standard deviation of the returns $X_{t,i}$, $i=1,\dots,n_{t}$.
To estimate the bandwidth for the nonparametric regression operator, we can use generalised cross-validation (GCV) to select the bandwidth automatically (see, e.g., Ferraty et al., 2012). The GCV aims to minimise an overall squared loss function between the estimated and observed responses among the data in the training sample.
@@ 3.2 Monte-Carlo simulation studies
    When generating the model $f_{t+1}=m(f_{t})\oplus\epsilon_{t}$, a first step is to be able to simulate $\epsilon_{0},\epsilon_{1},\ldots$, the probability density functions playing the role of model errors. By using the clr transformation, the density-on-density regression model is equivalent to $\mbox{clr}(f_{t+1})=\mbox{clr}\circ m(f_{t})+\mbox{clr}(\epsilon_{t})$. Instead of generating directly the errors $\epsilon_{t}$ in the Bayes Hilbert space $\mathcal{B}^{2}(I)$, it could be more convenient to simulated functional errors $\eta_{t}=\mbox{clr}(\epsilon_{t})$ in the corresponding Hilbert space $\mathcal{L}^{2}_{0}(I)$.
@@
[…]
@@
[…]
@@
    where $A_{tj}$’s are iid zero-mean real random variables (r.r.v.). In Figure 2, we present such model errors $\eta_{t}$ with $A_{t1}$’s,…,$A_{t5}$’s $\overset{iid}{\sim}$ $N(0,\sigma^{2})$.
@@
    Transform back to $\epsilon_{t}$’s with the inverse of the clr transformation: $\epsilon_{t}=\mbox{clr}^{-1}(\eta_{t})$. In Figure 3, we present 10 first model errors in Bayes space.
@@
    Let us define the regression operator $m:\,\mathcal{B}^{2}(I)\rightarrow\mathcal{B}^{2}(I)$ such that
[…]
    with $f$ and $g$ two given probability densities in $\mathcal{B}^{2}(I)$, $10<\rho_{0}<$ 1 and where $\underline{b}=\max[-1$, $\rho_{0}^{-1}(y+\rho_{0}-1)]$ and $\overline{b}=\min[1,\rho_{0}^{-1}(y-\rho_{0}+1)]$. Depending on the value of $\rho_{0}$, the function support may vary for the regression mean function. This regression operator has a nice interpretation in terms of convolution. Suppose that, for any $10<\rho_{0}<$ 1, $Z=\rho_{0}X+(1-\rho_{0})Y$ with $X$ and $Y$ two independent r.r.v. and let $f_{Z}$ (resp. $f_{X}$) be the PDF of $Z$ (resp. $X$). Then, we have $f_{Z}\,=\,m(f_{X})$ and $\text{Corr}(Z,\,X)=\rho_{0}$. The regression operator $m(.)$ corresponds to the transformation resulting from a convex combination of $X$ with another r.r.v. $Y$. Based on this remark, we propose the following scheme for simulating our time series of PDFs in the Bayes Hilbert space:
@@
    where the $\epsilon_{t}$’s are the model errors in Bayes space (built previously): 1) $m(f_{t})$represents the PDF of $X_{t+1}=\rho_{0}X_{t}+(1-\rho_{0})Y_{t+1}$ when $f_{t}$ is the PDF of $X_{t}$, 2) $f_{t+1}$is the obtained PDF when adding some model error $\epsilon_{t}$ to $m(f_{t})$.
@@
    Set $Y_{t}\,\sim\,g_{t}=TN(\mu_{t},\nu^{2})$ where $TN(\mu_{t},\nu^{2})$ is the truncated normal distribution over $[-1,1]$ derived from the normal distribution $N(\mu_{t},\nu^{2})$ where, for some constant $T>0$, $\mu_{t}=\cos(2\pi t/T)$.
@@
In Figure 4, we present a perspective plot of a simulated example with $150$ curves. We consider 201 grid points within a function support range $[-1,1]$. There are four tuning parameters to choose, namely $\sigma$ in step 2), $\rho_{0}$ in step 3), $\nu$ and $T$ in step 4). Let $\sigma=0.1$, $\rho_{0}=0.5$, $\nu=0.5$ and $T=150$.
@@
For one out of 100 replications, we generate a sample of densities under different tuning parameters. With the training samples, we produce one-step-ahead forecast densities using the nonparametric density-on-density regression with the Bayes NW estimator, the CoDa method, the LQDT, and functional principal component regression of Horta & Ziegelmann (2018). In Figure 5, we present one replication of holdout densities and their forecasts obtained by the five methods. For this example, we observe that the nonparametric density-on-density regression produces the best results, followed closely by the random walk. The Horta & Ziegelmann’s (2018) method suffers from the well-known tail problem when density forecasts can even be negative.
@@
To further assess the overall forecast accuracy, we repeat the simulation data generating process 100 times, each with different pseudo-random seeds. In Table 1, we present the one-step-ahead averaged KLDs between the actual holdout samples and their corresponding forecasts obtained from the five methods. Further, we present results using expanding window approach where the size of the training samples increases. As measured by the $\overline{\text{KLD}}$ criterion, the forecast accuracy relies heavily on the noise-to-signal ratio, characterised by the dependence parameter $\rho_{0}$ and the standard deviation $\sigma$ in the error term of the AR(1) structure. As the dependence parameter $\rho_{0}$ decreases from 0.75 to 0.25, we observe a decrease of $\overline{\text{KLD}}$ based on 100 replications. As the standard deviation in the error term of the AR(1) structure increases from 0.1 to 1, we also observe an increase of $\overline{\text{KLD}}$. As the sample size increases from $150$ to 550, we observe a general increase of $\overline{\text{nsr}}$ which does not translate to an increase of $\overline{\text{KLD}}$.
Table 1: One-step-ahead forecast accuracy of the nonparametric density-on-density regression with the Bayes NW estimator in the simulated data with 100 replications. We also compare the forecast accuracy with the CoDa, LQDT, HZ and RW methods, under three choices of $(\sigma,\rho_{0})$ with sample sizes of $T=150,350$ and $550$. The last 50 densities are the testing sample, while the remaining densities are the training sample. For the 100 replications, we compute the average noise-to-signal ratio, denoted by $\overline{\text{nsr}}$.
@@
We observe that the Horta & Ziegelmann’s (2018) method produces inferior accuracy because it ignores the density constraints. By obeying the density constraints, the remaining four methods improve the forecast accuracy. Among the four methods, the LQDT does not work well when functions have different finite supports, as evident from large $\overline{\text{KLD}}$ values. When the $\sigma$ value is smaller, the nonparametric density-on-density regression using the functional Bayes NW estimator is the chosen method with the smallest $\overline{\text{KLD}}$. When the $\sigma$ value is larger, and $\rho_{0}$ value is smaller, the CoDa method is preferred.
@@ 3.3 Analysis of a French COVID19 hospitalisation dataset
The French open data portal https://www.data.gouv.fr/fr/ allows the general public to access many data sources. Simple research on this site with the keyword COVID19 leads us to several data sets. In this subsection, we are interested in the dynamics of the pandemic in France at the departmental level and, more particularly, in the hospitalisation data. French territory is divided into 101 administrative areas named départements. In Figure 6, we plot 96 departments located inside Europe.
@@
The French COVID19 data, updated daily, are contained in a .csv file whose name starts with covid-hospit-2023 and can be downloaded https://www.data.gouv.fr/fr/datasets/donnees-hospitalieres-relatives-a-lepidemie-de-covid-19/. We are interested in studying the number of hospitalisations per 100,000 inhabitants for each department from March 18, 2020, to August 24, 2022. By focusing on cross-sectional data, we collect 96 daily hospitalisation ratios $(\mathcal{X}_{t,1},\mathcal{X}_{t,2},\dots,\mathcal{X}_{t,96})$ for each observed date $t=1,2,\dots,890$. By taking a simple average over days, we can single out the departments with minimum and maximum hospitalisation ratios. The department Vendée has the minimum hospitalisation ratio, with an average of 8.58 per 100,000 inhabitants. The department Territoire de Belfort has the maximum hospitalisation ratio, with an average of 50.52 per 100,000 inhabitants.
Since the number of hospitalisation is nonnegative, it is an unnecessary constraint that can be removed via a natural logarithm transformation, denoted by $\ln(\cdot)$. Let $Z_{t,i}=\ln(\mathcal{X}_{t,i}+c)$ with $c>0$ for $i=1,2,\dots,96$, where $c$ is an arbitrary small value such as $c=0.1$. Then, we apply a kernel density estimator to estimate the probability density function $f_{t}$ of $Z_{t}$ without positive constraint. Computationally, the estimation is done via density function in [Uncaptioned image] (R Core Team, 2025) using Silverman’s ROT bandwidth selection and truncated Gaussian kernel function. In Figure 7, we present the 890 estimated densities.
@@
In Figure 8, we present the 296 one-step-ahead density forecasts of the four methods considered. Since the regression mean function (i.e., signal) is unknown, the random walk is not considered a benchmark method. Compared with the holdout densities, we found the nonparametric density-on-density regression with the Bayes NW estimator performs the best. Similar to the CoDa method, it also uses the log-ratio transformation to remove constraints. A univariate functional time series forecasting method or a nonparametric function-on-function is then applied to obtain the forecasts. The forecasts are then back-transformed via the inverse log-ratio transformation. The nonparametric density-on-density regression with the Bayes NW estimator generally produces smoother density forecasts than the ones obtained from the CoDa method. The intuition is that the former one is a weighted average of the past densities and tends to be smooth.
@@
Table 2: From the 296 KLDs, we present summary statistics for the four methods.
          Statistic           Bayes NW           CoDa           LQDT           HZ
          Min.           0.0041           0.0051           0.0171           0.0176
          1st Qu.           0.0188           0.0197           1.0127           0.0727
          Median           0.0293           0.0324           4.1861           0.1747
          Mean           0.0420           0.0476           8.4794           0.2014
          3rd Qu.           0.0524           0.0554           12.8809           0.2910
          Max.           0.5241           0.4243           33.2268           0.9041
@@ 3.4 Analysis of age-specific life-table death counts in the United States
While it requires a way of estimating densities using a kernel estimator in Section 3.3, sometimes, we observe a time series of densities, such as period life table in demography and actuarial studies. A period life table is a table which shows, for each age, the probability that a person of that age will die before the next birthday. In the first age group, the initial number of alive is 100,000, while the remaining number of alive is 0 in the last age group. The sum of life-table death count is 100,000 every year, and the values are nonnegative, so these observations can be viewed as densities normalised to 100,000 rather than 1. By modelling the life-table death counts, we could understand a redistribution of survival probabilities, where deaths at younger ages gradually shift towards older ages. In Figure 1, we present rainbow plots of the female and male age distributions of life-table death counts in the USA from 1933 to 2023 in a single-year group.
@@
Figure 9: Holdout and forecast female and male life-table death counts from 2001 to 2023. The forecast life-table death counts were obtained via the nonparametric density-on-density regression with Bayes NW estimator, CoDa method, LQDT method, and HZ method, where the optimal bandwidth parameter was selected by the GCV.
@@
Figure 10: Forecast accuracy (MSPE ratios) between the nonparametric density-on-density regression using the Bayes NW, CoDa method, LQD method, and HZ method for forecasting one-step-ahead age-specific life-table death counts from 2001 to 2023 in the USA.
We evaluate and compare the KLD among the Bayes NW, CoDa, LQDT and HZ methods. The KLD of the nonparametric density-on-density regression with the Bayes NW estimator is the smallest among the four methods considered. The inferior performance of the LQDT method is primarily due to the left boundary problem, where there is a comparably higher number of deaths at the infant age.

### v2

@@ 3.1 Implementation
Since actual densities in practical applications are not often observable, we work with estimated densities $\widehat{f}_{1},\ldots,\widehat{f}_{N+1}$ which are outputs of the kernel density estimator defined with (5) where $K_{\text{kde}}(\cdot)$ stands for a kernel function and $h_{t}$ denotes a bandwidth for period $t$. We consider the truncated Gaussian kernel.
In our nonparametric density-on-density regression, we consider how to select the optimal bandwidths for estimating the nonparametric regression operator in (1) and for estimating densities in (5). To estimate densities from observed data, we use a kernel density estimator with the bandwidth selected by Silverman’s rule-of-thumb (ROT) (see also Kokoszka et al., 2019). Kokoszka et al. (2019) used Silverman’s ROT to select the bandwidth, which leads to $\widehat{h}_{t}=2.34\times\widehat{\sigma}_{t}\times n^{-1/5}$, where $\widehat{\sigma_{t}}$ denotes the sample standard deviation of $X_{t,i}$, $i=1,\dots,n$.
To estimate the bandwidth for the nonparametric regression operator, we can use generalised cross-validation (GCV) to select the bandwidth automatically (see, e.g., Ferraty et al., 2012a). The GCV aims to minimise an overall squared loss function between the estimated and observed responses among the data in the training sample.
@@ 3.2 Monte-Carlo simulation studies
    When generating the model $f_{t+1}=m(f_{t})\oplus\epsilon_{t}$, a first step is to be able to simulate $\epsilon_{0},\epsilon_{1},\ldots$, the probability density functions playing the role of model errors. By using the clr transformation, the density-on-density regression model is equivalent to $\mbox{clr}(f_{t+1})=\mbox{clr}\circ m(f_{t})+\mbox{clr}(\epsilon_{t})$. Instead of generating directly the errors $\epsilon_{t}$ in the Bayes Hilbert space $\mathcal{B}_{2}(I)$, it could be more convenient to simulate functional errors $\eta_{t}=\mbox{clr}(\epsilon_{t})$ in the corresponding Hilbert space $\mathcal{L}^{2}_{0}(I)$.
@@
    According to the properties of the clr transformation and the zero-mean constraint in $\mathcal{B}_{2}(I)$ of $\epsilon_{t}$ given $f_{t}$, the model error $\eta_{t}$ has to fulfill two constraints: $\text{E}(\eta_{t}|f_{t})=0$, and for all $u\in$ II, $\int_{I}\eta_{t}(u)\,du=0$. A way to generate such $\eta_{t}$’s is to use the trigonometric functions $\phi_{1},\phi_{2},\ldots$ defined on I=[1, 1]I=[-1,\,1]: $\forall k=1,2\ldots,\ \phi_{2k-1}(u)=\cos(k\pi\,u)$ and $\phi_{2k}(u)=\sin(k\pi\,u)$. One can remark that the $\phi_{k}$’s integrate to 0 for any non-null integer $k$. For any $u\in[-1,\,1]$, and any subset $\mathcal{K}$ of nonnull integers, set
@@
    If $\text{E}(A_{tk}|f_{t})=0$, then $\text{E}(\eta_{t}|f_{t})=0$ and the integral property of the Fourier basis elements results in $\int_{-1}^{1}\eta_{t}(u)\,du=0$. For instance, one can set
@@
    where $A_{tj}$’s are iid zero-mean rrv. In Figure 2, we present such model errors $\eta_{t}$ with $A_{t1}$$A_{t5}’s$ $\overset{iid}{\sim}$ $N(0,\sigma^{2})$.
@@
    Transform back to $\epsilon_{t}$’s with the inverse of the clr transformation: $\epsilon_{t}=\mbox{clr}^{-1}(\eta_{t})$. In Figure 3, we present 10 first model errors in the Bayes space.
@@
    Let us define the regression operator $m:\,\mathcal{B}_{2}(I)\rightarrow\mathcal{B}_{2}(I)$ such that
    $m(f)(y)\,=\,(1-\rho_{0})^{-1}\left\{\int_{\underline{b}}^{\overline{b}}f(x)\,g\left(\frac{y-\rho_{0}\,x}{1-\rho_{0}}\right)\,dx\right\}1_{[-1,\,1]}(y)$
    with $f$ and $g$ two given probability densities in $\mathcal{B}_{2}(I)$, $0<\rho_{0}<1$ and where $\underline{b}=\max[-1$, $\rho_{0}^{-1}(y+\rho_{0}-1)]$ and $\overline{b}=\min[1,\rho_{0}^{-1}(y-\rho_{0}+1)]$. Depending on the value of $\rho_{0}$, the function support may vary for the regression mean function. This regression operator has a nice interpretation in terms of convolution. Suppose that for any $0<\rho_{0}<1$, $Z=\rho_{0}X+(1-\rho_{0})Y$ with $X$ and $Y$ two independent rrv and let $f_{Z}$ (resp. $f_{X}$) be the PDF of $Z$ (resp. $X$). Then, we have $f_{Z}\,=\,m(f_{X})$ and $\text{Corr}(Z,\,X)=\rho_{0}$. The regression operator $m(.)$ corresponds to the transformation resulting from a convex combination of $X$ with another rrv $Y$. Based on this remark, we propose the following scheme for simulating our time series of PDFs in the Bayes Hilbert space:
@@
    where the $\epsilon_{t}$’s are the model errors in the Bayes space (built previously): 1) $m(f_{t})$represents the PDF of $X_{t+1}=\rho_{0}X_{t}+(1-\rho_{0})Y_{t+1}$ when $f_{t}$ is the PDF of $X_{t}$, 2) $f_{t+1}$is the obtained PDF when adding some model error $\epsilon_{t}$ to $m(f_{t})$.
@@
    Set $Y_{t}\,\sim\,g_{t}=TN(\mu_{t},\nu^{2})$ where $TN(\mu_{t},\nu^{2})$ is the truncated normal distribution over [1,1][-1,1] derived from the normal distribution $N(\mu_{t},\nu^{2})$ where, for some constant $T>0$, $\mu_{t}=\cos(2\pi t/T)$.
@@
In Figure 4, we present a perspective plot of a simulated example with $150$ curves. We consider 201 grid points within a function support range [1,1][-1,1]. There are four tuning parameters to choose, namely $\sigma$ in step 2), $\rho_{0}$ in step 3), $\nu$ and $T$ in step 4). Let $\sigma=0.1$, $\rho_{0}=0.5$, $\nu=0.5$ and $T=150$.
@@
For one out of 100 replications, we generate a sample of densities under different tuning parameters. With the training samples, we produce one-step-ahead forecast densities using the nonparametric density-on-density regression with the Bayes NW estimator, the CoDa method, the LQDT, functional principal component regression of Horta & Ziegelmann (2018), and random walk. In Figure 5, we present one replication of holdout densities and their forecasts obtained by the five methods. For this example, we observe that the nonparametric density-on-density regression produces the best results, followed closely by the random walk. The functional principal component regression of Horta & Ziegelmann (2018) suffers from the well-known tail problem when density forecasts can even be negative.
@@
To further assess the overall forecast accuracy, we repeat the simulation data generating process 100 times, each with different pseudo-random seeds. In Table 1, we present the one-step-ahead averaged KLDs between the actual holdout samples and their corresponding forecasts obtained from the five methods. In addition, we present results using an expanding window approach in which the size of the training samples increases. As measured by the $\overline{\text{KLD}}$ criterion, the forecast accuracy relies heavily on the noise-to-signal ratio, characterised by the dependence parameter $\rho_{0}$ and the standard deviation $\sigma$ in the error term of the AR(1) structure. As the dependence parameter $\rho_{0}$ decreases from 0.75 to 0.25, we observe a decrease of $\overline{\text{KLD}}$ based on 100 replications. As the standard deviation in the error term of the AR(1) structure increases from 0.1 to 1, we also observe an increase of $\overline{\text{KLD}}$. As the sample size increases from $150$ to 550, we observe a general increase of $\overline{\text{nsr}}$ which does not translate to an increase of $\overline{\text{KLD}}$.
Table 1: One-step-ahead forecast accuracy of the nonparametric density-on-density regression, simulated data with 100 replications. We consider three choices of $(\sigma,\rho_{0})$ with sample sizes of $T=150,350$ and $550$. The last 50 densities are the testing sample, while the remaining densities are the training sample. For the 100 replications, we compute the average noise-to-signal ratio, denoted by $\overline{\text{nsr}}$. HZ stands for Horta & Ziegelmann’s (2018) method.
@@
We observe that the Horta & Ziegelmann’s (2018) method produces inferior accuracy because it ignores the density constraints. By obeying the density constraints, the remaining four methods improve the forecast accuracy. Among the four methods, the LQDT does not work well when functions have different finite supports, as is evident from the large $\overline{\text{KLD}}$ values. When the value of $\sigma$ is lower, the nonparametric density-on-density regression using the functional Bayes NW estimator is the method chosen with the smallest $\overline{\text{KLD}}$. When the value of $\sigma$ is higher and the value of $\rho_{0}$ is lower, the CoDa method is preferred.
@@ 3.3 Analysis of a French COVID19 hospitalisation dataset
The French open data portal https://www.data.gouv.fr/fr/ allows the general public to access many data sources. Simple research on this site with the keyword COVID19 leads us to several data sets. In this subsection, we are interested in the dynamics of the pandemic in France at the departmental level and, more particularly, in the hospitalisation data. The French territory is divided into 101 administrative areas named départements. In Figure 6, we plot 96 departments located within Europe.
@@
The French COVID19 data, updated daily, are contained in a .csv file whose name starts with covid-hospit-2023 and can be downloaded https://www.data.gouv.fr/fr/datasets/donnees-hospitalieres-relatives-a-lepidemie-de-covid-19/. We are interested in studying the number of hospitalisations per 100,000 inhabitants for each department from March 18, 2020, to August 24, 2022. Using cross-sectional data, we collect 96 daily hospitalisation ratios $(\mathcal{X}_{t,1},\mathcal{X}_{t,2},\dots,\mathcal{X}_{t,96})$ for each observed date $t=1,2,\dots,890$. By taking a simple average over days, we can identify departments with minimum and maximum hospitalisation ratios. The department Vendée has the minimum hospitalisation ratio, with an average of 8.58 per 100,000 inhabitants. The Territoire de Belfort department has the highest hospitalisation ratio, with an average of 50.52 per 100,000 inhabitants.
Since the number of hospitalisation is nonnegative, the nonnegativity constraint can be removed by applying the natural logarithm transformation, denoted by $\ln(\cdot)$. Let $X_{t,i}=\ln(\mathcal{X}_{t,i}+c)$ with $c>0$ for $i=1,2,\dots,96$, where $c$ is an arbitrary small value such as $c=0.1$. Then, we apply a kernel density estimator to estimate the probability density function $f_{t}$ based on the sample $\{X_{t,1},\ldots,X_{t,96}\}$ without a positive constraint. Computationally, the estimation is implemented via the density function in [Uncaptioned image] (R Core Team, 2025) using Silverman’s ROT bandwidth selection and the truncated Gaussian kernel function. In Figure 7, we present the 890 estimated densities.
@@
In Figure 8, we present the 296 one-step-ahead density forecasts of the four methods considered. Since the regression mean function (i.e., signal) is unknown, the random walk is not considered a benchmark method. Compared with the holdout densities, we found that nonparametric density-on-density regression with the Bayes NW estimator performs the best. Similar to the CoDa method, it also uses the log-ratio transformation to remove constraints. A univariate functional time series forecasting method or a nonparametric function-on-function is then applied to obtain the forecasts. The forecasts are then back-transformed via the inverse log-ratio transformation. The nonparametric density-on-density regression with the Bayes NW estimator generally produces smoother density forecasts than the ones obtained from the CoDa method. The intuition is that the former is a weighted average of the past densities and tends to be smooth.
@@
Among the methods, the LQDT exhibits a different pattern. Using other time-series forecasting methods, they do not improve the forecast accuracy measured by the KLD. We suspect that the issue may lie in the chosen reference point. Without an adequate reference point, the LQDT would only be defined up to an arbitrary constant.
@@
Table 2: From the 296 KLDs, we present summary statistics for the four methods, as well as the random-walk (RW) method. HZ stands for Horta & Ziegelmann’s (2018) method.
   Statistic    Bayes NW    CoDa    LQDT    HZ    RW
   Min.    0.0041    0.0051    0.0171    0.0176    0.0008
   1st Qu.    0.0188    0.0197    1.0127    0.0727    0.0185
   Median    0.0293    0.0324    4.1861    0.1747    0.0331
   Mean    0.0420    0.0476    8.4794    0.2014    0.0450
   3rd Qu.    0.0524    0.0554    12.8809    0.2910    0.0561
   Max.    0.5241    0.4243    33.2268    0.9041    0.4461
@@ 3.4 Analysis of age-specific life-table death counts in the United States
While it requires a way of estimating densities using a kernel estimator in Section 3.3, sometimes, we observe a time series of densities, such as period life table in demography and actuarial studies. A period life table is a table that shows, for each age, the probability that a person of that age will die before the next birthday. In the first age group, the initial number of alive is 100,000, while the remaining number of alive is 0 in the last age group. The sum of life-table death count is 100,000 every year, and the values are nonnegative, so these observations can be viewed as densities normalised to 100,000 rather than 1. By modelling the life-table death counts, we could understand a redistribution of survival probabilities, where deaths at younger ages gradually shift towards older ages. In Figure 1, we present rainbow plots of the female and male age distributions of life-table death counts in the U.S. from 1933 to 2023 in a single-year group.
@@
Figure 9: Holdout and forecast female and male life-table death counts from 2001 to 2023. The forecast life-table death counts were obtained via the nonparametric density-on-density regression with Bayes NW estimator, CoDa method, LQDT method, and functional principal component regression of Horta & Ziegelmann (2018), where the optimal bandwidth parameter was selected by the GCV.
@@
Figure 10: Forecast accuracy (MSPE ratios) between the nonparametric density-on-density regression using the Bayes NW, CoDa method, LQD method, and functional principal component regression of Horta & Ziegelmann (2018) for forecasting one-step-ahead age-specific life-table death counts from 2001 to 2023 in the USA.
We evaluate and compare the KLD among the Bayes NW, CoDa, LQDT methods, and functional principal component regression of Horta & Ziegelmann (2018). The KLD of the nonparametric density-on-density regression with the Bayes NW estimator is the smallest among the four methods considered. The inferior performance of the LQDT method is primarily due to the left boundary problem, where there is a comparably higher number of deaths at the infant age.